# A relativistic model of b-EMRI systems and their gravitational radiation

### v1

@@ II.4 Teukolsky formalism
where $s=-2$ is the spin weight corresponding to a gravitational field. The master function $\psi$ is related to the Weyl scalar by $\psi=\rho^{-4}\psi_{4}$, where $\psi_{4}=-C_{\alpha\beta\gamma\delta}n^{\alpha}\bar{m}^{\beta}n^{\gamma}\bar{m}^{\delta}$ is the Weyl scalar given by the Newman-Penrose (NP) formalism [94]. The term $\mathcal{T}$ describes the source of perturbation.
@@
The solution to the inhomogeneous radial equation can be constructed from the homogeneous solutions using the Green function method, which results in
@@
The expressions for $A_{nn0}$, $A_{\bar{m}n0}$, $A_{\bar{m}\bar{m}0}$, $A_{\bar{m}n1}$, $A_{\bar{m}\bar{m}1}$, and $A_{\bar{m}\bar{m}2}$ can be found in Appendix B. Note that here we have assumed the stress-energy of two free monopole particles. We discuss this assumption in further detail in Section V.3.
Now we analyze the asymptotic behavior of Eq. (17). When $r\to r_{+}$, the lower and upper limits in the integration of $Z^{\mathrm{in}}_{lm}(r,\omega)$ coincide, leading to $Z^{\mathrm{in}}_{lm}(r,\omega)=0$. For the same reason, when $r\to\infty$, $Z^{\mathrm{up}}_{lm}(r,\omega)=0$. For our b-EMRI, the distance to the source $r$ is much greater than the size of the system, $p$. Therefore, effectively we have $r/p\to\infty$, and in this case the only non-vanishing term is
@@ II.5 Expansion of the source term
$Z^{\mathrm{in}(0)}_{lm}(\omega)=$ $\frac{\mu}{\mathcal{W}_{lm}(\omega)}\int_{-\infty}^{\infty}Z_{lm}^{(0)}(\omega)\ \mathrm{d}\tau$ (33)
@@
$=$ $\frac{2\pi\mu}{u^{t}\mathcal{W}_{lm}(\omega)}\delta\left(\omega-m\frac{u^{\varphi}}{u^{t}}\right)\mathcal{Z}_{lm}^{(0)}(r_{0},\theta_{0},\omega)$
$=$ $\frac{2\pi\mu}{u^{t}\mathcal{W}_{lm}(\omega_{m}^{(0)})}\mathcal{Z}^{(0)}_{lm}(r_{0},\theta_{0},\omega_{m}^{(0)}).$
@@
For the first-order waveform, we find that the frequencies satisfy the condition
$\omega_{mwv}^{(1)}=m\frac{u^{\varphi}}{u^{t}}+w\frac{\Omega}{u^{t}}+v\frac{\omega_{IB}}{u^{t}},$ (36)
where $m=-l,\dots,l$, $w=0,\pm 1$, and $v=\pm 1$. Therefore, we have six split first-order modes for each $m$. The corresponding amplitudes are
@@
where $Z^{(1)}_{lmwv}\equiv Z^{(1)}_{lm}(r_{0},\theta_{0},\omega_{mwv}^{(1)})$ is the split first-order amplitude and $\mathcal{W}_{lmwv}\equiv\mathcal{W}_{lm}(\omega_{mwv})$ is the Wronskian.
@@
$\omega_{mwv}^{(2)}=m\frac{u^{\varphi}}{u^{t}}+w\frac{\Omega}{u^{t}}+v\frac{\omega_{IB}}{u^{t}},$ (38)
@@ II.7 Adiabatic evolution of the system
Integrating the $03$ component over the solid angle yields the average energy fluxes at infinity and at the event horizon $r_{+}$. They can be expressed as
@@
where $q_{1}$and $q_{2}$ are functions of $\mathcal{E}$, $\mathcal{L}_{z}$ and $p$, and these two functions are derived as follows. Since circular orbit remains circular during radiation reaction [96, 97, 98, 99], we have $\dot{R}=0$ and $\dot{R}^{\prime}=0$ according to the radial geodesic equation defined in Eq. (5). The former equation relates the energy and angular-momentum fluxes to the properties of the circular-equatorial orbits, which is
@@ III.1 Frequencies and amplitudes of the modes
$\Omega_{G}\equiv\frac{u^{\varphi}}{u^{t}},\quad\Omega_{P}\equiv\frac{\Omega}{u^{t}},\quad\Omega_{I}\equiv\frac{\omega_{IB}}{u^{t}}.$ (53)
@@ IV.1 Amplitude, Flux, and resonant excitation of QNMs
Table 2: The zeroth and second order amplitudes and their fractions of main modes, with parameters $a=0.9M$, $p=10M$, $\epsilon=10^{-5}$, $d=4000\mu$, $\tilde{\psi}=\pi/4$, and $\tilde{\theta}=\pi/3$.
@@
It is remarkable that the ratios $A_{lmwv}$ for the high-frequency modes with $v=\pm 2$ can approach unity in some cases. These relative second-order amplitudes are about $\epsilon^{-1}$ times larger than those of the low-frequency modes. Moreover, for each $m$, there are ten high-frequency modes corresponding to $w=0,\pm 1,\pm 2,\ v=\pm 2$. Therefore, the fluxes are amplified by another order of magnitude. Such a large flux due to the high-frequency modes should cause the b-EMRI system to evolve faster than a standard EMRI. We notice that an earlier study using the MPD equation to solve the evolution of b-EMRI also revealed the appearance of two high-frequency modes (see Eq. (30) in Ref. [65]). However, the authors disregarded these two terms in the calculation of the flux, which potentially could result in a slower evolution of their b-EMRI.
@@
From the table we find that that the contributions from the b-EMRI high-frequency modes are substantial, particularly during the early inspiral stage. This leads to a faster evolution of the b-EMRI system compared to a standard EMRI. In contrast, the low-frequency modes contribute only at the $\mathcal{O}(\epsilon)$ order. The behavior of the angular momentum fluxes mirrors that of the energy fluxes and hence are not shown here.
@@
Figure 4: Waveform snapshots with different setting of $d$. Other parameters are fixed at $a=0.9M$, $p=10M$, $\epsilon=10^{-5}$, $\tilde{\psi}=\pi/4$, $\tilde{\theta}=\pi/3$ with viewing angles $\theta=\pi/3$ and $\varphi=\pi/6$.
@@ IV.2 Adiabatic trajectory and waveform
In this subsection, we study the adiabatic evolution of the b-EMRI system with different configurations and compare it with a standard EMRI. We continue to use the parameters $a=0.9M$, $p=10M$, $\epsilon=10^{-5}$, $d=4000\mu$, $\tilde{\psi}=\pi/4$, and $\tilde{\theta}=\pi/3$. We also adopt the configurations defined in Eq. (60). Eq. (60a) corresponds to the evolution of a standard EMRI. Eq. (60b) represents the evolution of a b-EMRI driven only by the low-frequency modes, and we refer to this case as “Configuration 1”. Eq. (60c) includes all modes, which we refer to as “Configuration 2”.
@@
These results suggest that the high-frequency modes contribute significantly more to the evolution of the b-EMRI system than the low-frequency modes. As a result, the b-EMRI system evolve much faster than a standard EMRI in the early inspiral phase. This observation extends beyond the current understanding of b-EMRI system evolution. For example, in Fig. 2 of Ref. [65] where the contribution of the high-frequency modes are not included, we see a much slower evolution of the b-EMRI system which is consistent with what we have derived in Fig. 6(a).
@@
Table 4: Energy fluxes due to different modes when $a=0.9M$, $\epsilon=10^{-5}$, $\tilde{\psi}=\pi/4$, and $\tilde{\theta}=\pi/3$
@@ V.1 Comparisons with previous works
Later, Ref. [65] treats the inner binary as a mass distribution and expands it into multipole moments to compute the source term. To evolve the outer orbit, the authors treat the BBH as a single spinning body and use the MPD equation to compute its orbit around the SMBH. This approach ensures conservation of the stress-energy tensor and inspires us to discuss the validity of our model in Sec. V.3. However, in their study the authors omitted the high-frequency source terms because the frequencies are outside the LISA band. Our results, nevertheless, shows that these dropped high-frequency modes, if resonate with the central SMBH’s QNMs, will enhance energy dissipation.
Notably, the IB is allowed to evolve and merger in Ref. [65], so the b-EMRI will be an important multi-band GW source which can produce both high-frequency GW in the LIGO/Virgo band and low-frequency GW in the LISA band [40]. The merger of the IB also leads to a sudden mass loss making the low-frequency waveform more distinguishable. These potential multi-band observational aspects are not discussed in our work.
@@ V.2 Parameter space
Throughout this work, we assume no evolution of the IB, which means that the IB is neither tidally disrupted nor excited to high eccentricities. The stability of a hierarchical three-body system is a classical problem [110]. The maximum distance for the two bodies in the IB to remain stable is determined by the Hill radius, $R_{H}:=(\epsilon/3)^{1/3}p$, where Newtonian gravity is assumed. Since here we are working in the strong-gravity regime, we require $d<R_{H}/3$, to be conservative. Given our fiducial model in which $p=10M$ and $\epsilon=10^{-5}$, we find $d<0.05M$. In other words, an IB with $d=4000\mu=0.04M$ is stable in this case. Additionally, since we assume Newtonian motion for the inner orbital motion, $d$ should be sufficiently large to avoid strong relativistic correction. For this latter reason, we set $d>1000\mu=1000M\epsilon$.
Another important assumption in this work is that the inner binary follows a circular orbit. However, dynamical effects in triple systems, such as the Kozai-Lidov mechanism [111, 112], could excite the eccentricity of the IB if the inclination angle $\tilde{\theta}$ between the inner and outer orbits exceeds a critical value. The timescale of the eccentric Kozai-Lidov mechanism (EKM) was derived in Ref. [113, 114], which gives
@@
Figure 7: Parameter space of a b-EMRI. The regions below the solid (dashed) lines satisfy the EKM (stability) condition when $\tilde{p}=20,10,5$. The grey shaded region represents a IB in the mass range of $10M_{\odot}<\mu<100M_{\odot}$, where we have assumed $M=4\times 10^{6}M_{\odot}$. The red star denotes our fiducial model in which $\epsilon=10^{-5},\ \tilde{d}=4000$.
@@ V.3 Validity of the model
In Sec. II.7 we demonstrated how to track the evolution of the system ‘quasi-adiabaticly’. In particular, the evolution of the parameter $p$ is derived assuming osculating geodesic motion of the OB. In this work, however, we follow Ref. [63, 64] and write the stress-energy tensor of the IB as
@@
While we have not yet included this interaction term in the stress-energy when solving the Teukolsky equation, we remark that it is straight forward to add in follow up work as a linear correction to the Teukolsky amplitudes (and therefore fluxes). Without accounting for that term, the waveform model should not be used in parameter estimation studies for example. However, its neglect does not change the qualitative features of our findings.
@@ Acknowledgements.
This work is supported by the National Key Research and Development Program of China (Grant No. 2021YFC2203002) and the National Natural Science Foundation of China (Grant No. 11991053). Y.Y. acknowledges NUS Physics Department for providing comfortable office environment and warm care. Y.Y. would like to thank Haolan Zheng and Han Yan for several inspirational discussions. The authors thank Soichiro Isoyama, Wenbiao Han, and Xilong Fan for insightful feedback on a draft of this paper. J.M. acknowledges support by the NUS Faculty of Science, under the research grant 22-5478-A0001.
@@ Appendix A Transformation from the LIF to BL coordinates
$c_{01}=$ $\frac{\mathcal{M}\mathcal{M}_{\varphi}+\gamma_{\varphi\varphi}^{1/2}\nu}{4\gamma_{\varphi\varphi}^{1/2}\mathcal{M}/\gamma}\Big{\{}(1-\sin\tilde{\theta})\sin\left[\left(\Omega+\omega_{IB}\right)\tau+\tilde{\psi}\right]$
$+(1+\sin\tilde{\theta})\sin\left[\left(\Omega-\omega_{IB}\right)\tau+\tilde{\psi}\right]\Big{\}},$ (72a)
$c_{11}=$ $\frac{1}{4\gamma_{rr}^{1/2}}\Big{\{}(1-\sin\tilde{\theta})\cos\left[\left(\Omega+\omega_{IB}\right)\tau+\tilde{\psi}\right]$
$+(1+\sin\tilde{\theta})\cos\left[\left(\Omega-\omega_{IB}\right)\tau+\tilde{\psi}\right]\Big{\}},$ (72b)
$c_{21}=$ $\frac{1}{2\gamma_{\theta\theta}^{1/2}}\cos\tilde{\theta}\sin\left(\omega_{IB}\tau\right),$ (72c)
@@
$c_{31}=$ $\frac{\gamma}{4\gamma_{\varphi\varphi}^{1/2}}\Big{\{}(1-\sin\tilde{\theta})\sin\left[\left(\Omega+\omega_{IB}\right)\tau+\tilde{\psi}\right]$
$+(1+\sin\tilde{\theta})\sin\left[\left(\Omega-\omega_{IB}\right)\tau+\tilde{\psi}\right]\Big{\}},$ (72d)
$c_{02}=$ $\frac{(\mathcal{M}\mathcal{M}_{\varphi}+\gamma_{\varphi\varphi}^{1/2}\nu)\kappa_{\varphi}-(\gamma_{\varphi\varphi}^{1/2}+\mathcal{M}\mathcal{M}_{\varphi}\nu)\Omega}{32\gamma_{\varphi\varphi}^{1/2}\mathcal{M}/\gamma}$
×$\times\Big{\{}\left(1-\sin\tilde{\theta}\right)^{2}\sin\left[2\left(\Omega+\omega_{IB}\right)\tau+2\tilde{\psi}\right]$
$+\cos^{2}\tilde{\theta}\sin\left(2\Omega\tau+2\tilde{\psi}\right)$
$+\left(1+\sin\tilde{\theta}\right)^{2}\sin\left[2\left(\Omega-\omega_{IB}\right)\tau+2\tilde{\psi}\right]\Big{\}},$ (72e)
$c_{12}=$ $\frac{\kappa_{r}+\kappa_{\varphi}}{64\gamma_{rr}^{1/2}}\left(1-\sin\tilde{\theta}\right)^{2}\cos\left[2\left(\Omega+\omega_{IB}\right)\tau+2\tilde{\psi}\right]$
$+\frac{\kappa_{r}+\kappa_{\varphi}}{64\gamma_{rr}^{1/2}}\left(1+\sin\tilde{\theta}\right)^{2}\cos\left[2\left(\Omega-\omega_{IB}\right)\tau+2\tilde{\psi}\right]$
$+\frac{\kappa_{r}+\kappa_{\varphi}}{64\gamma_{rr}^{1/2}}\cos^{2}\tilde{\theta}\cos\left(2\Omega\tau+2\tilde{\psi}\right)$
$+\frac{\kappa_{r}+4\kappa_{\theta}-\kappa_{\varphi}}{64\gamma_{rr}^{1/2}}\cos^{2}\tilde{\theta}\cos\left(2\omega_{IB}\tau\right)$
$+\frac{\kappa_{r}-\kappa_{\varphi}}{32\gamma_{rr}^{1/2}}\left(1+\sin^{2}\tilde{\theta}\right)-\frac{\kappa_{\theta}}{16\gamma_{rr}^{1/2}}\cos^{2}\tilde{\theta},$ (72f)
@@
$c_{22}=$ $\frac{\kappa_{\theta}\cos\tilde{\theta}}{16\gamma_{\theta\theta}^{1/2}}\Big{\{}\left(1-\sin\tilde{\theta}\right)\sin\left[\left(\Omega+2\omega_{IB}\right)\tau+\tilde{\psi}\right]$
$-\left(1+\sin\tilde{\theta}\right)\sin\left[\left(\Omega-2\omega_{IB}\right)\tau+\tilde{\psi}\right]$
$+2\sin\tilde{\theta}\sin\left(\Omega\tau+\tilde{\psi}\right)\Big{\}},$ (72g)
@@
$c_{32}=$ $\frac{\kappa_{\varphi}-\nu\Omega}{32\gamma_{\varphi\varphi}^{1/2}/\gamma}\times\Big{\{}\left(1-\sin\tilde{\theta}\right)^{2}\sin\left[2\left(\Omega+\omega_{IB}\right)\tau+2\tilde{\psi}\right]$
@@
$+\cos^{2}\tilde{\theta}\sin\left(2\Omega\tau+2\tilde{\psi}\right)$
@@
$+\left(1+\sin\tilde{\theta}\right)^{2}\sin\left[2\left(\Omega-\omega_{IB}\right)\tau+2\tilde{\psi}\right]\Big{\}}.$ (72h)
As we can see, these coefficients are all function of $\tau$.

### v2

@@ II.4 Teukolsky formalism
where $s=-2$ is the spin weight corresponding to a gravitational field. The master function $\psi$ is related to the Weyl scalar by $\psi=\rho^{-4}\psi_{4}$, where $\rho=(r-ia\cos\theta)^{-1}$ and $\psi_{4}=-C_{\alpha\beta\gamma\delta}n^{\alpha}\bar{m}^{\beta}n^{\gamma}\bar{m}^{\delta}$ is the Weyl scalar given by the Newman-Penrose (NP) formalism [95]. The source term $\hat{T}=2\rho^{-4}T_{4}$ contains information about the two small BHs, which act as the perturbers. The expression for $T_{4}$ can be found in Ref. [94].
@@
The solution to the inhomogeneous radial equation can be constructed from the homogeneous solutions using the Green’s function method, which results in
@@
The expressions for $A_{nn0}$, $A_{\bar{m}n0}$, $A_{\bar{m}\bar{m}0}$, $A_{\bar{m}n1}$, $A_{\bar{m}\bar{m}1}$, and $A_{\bar{m}\bar{m}2}$ can be found in Appendix B.
Here we emphasize that we have derived the source term in Eq. (22) by modeling the IB’s stress-energy as that of two free monopole particles, similar to the approach used in Ref. [65], for example. To this end, the model in this work neglects the stress-energy due to the interaction between the two particles in the IB, which would contribute additional corrections to Eq. (22). We further elucidate this point in the discussion in Sec. V.3 and construct the relevant interactive stress-energy in Appendix D. We remind readers that our new approach to modeling b-EMRI systems is sufficient to draw qualitative conclusions, but still needs improvement for high-precision parameter estimation.
Now we analyze the asymptotic behavior of Eq. (17). When $r\to r_{+}$, the lower and upper limits in the integration of $Z^{\mathrm{in}}_{lm}(r,\omega)$ coincide, leading to $Z^{\mathrm{in}}_{lm}(r,\omega)=0$. For the same reason, when $r\to\infty$, $Z^{\mathrm{up}}_{lm}(r,\omega)=0$. For an observer in the solar system, the distance to the source $r$ is much larger than the size of the b-EMRI system, $p$. Therefore, we have $r/p\to\infty$, and in this case what is detectable by us is
@@ II.5 Expansion of the source term
$Z^{\mathrm{in}(0)}_{lm}(\omega)=$ $\frac{\mu}{\mathcal{W}_{lm}(\omega)}\int_{-\infty}^{\infty}Z_{lm}^{(0)}(\tau,\omega)\ \mathrm{d}\tau$ (33)
@@
$=$ $\frac{2\pi\mu}{u^{t}\mathcal{W}_{lm}(\omega)}\delta\left(\omega-m\frac{u^{\varphi}}{u^{t}}\right)\mathcal{Z}_{lm}^{(0)}(r_{0},\theta_{0},\omega).$
@@
For the first order waveform, we find that the frequencies satisfy the condition
$\omega_{mwv}^{(1)}=m\frac{u^{\varphi}}{u^{t}}+w\frac{\Omega}{u^{t}}+v\frac{\omega_{\mathrm{IB}}}{u^{t}},$ (36)
where $m=-l,\dots,l$, $w=0,\pm 1$, and $v=\pm 1$. Therefore, we have six split first order modes for each $m$. The corresponding amplitudes are
@@
where $Z^{(1)}_{lmwv}\equiv Z^{(1)}_{lm}(r_{0},\theta_{0},\omega_{mwv}^{(1)})$ is the split first order amplitude and $\mathcal{W}_{lmwv}\equiv\mathcal{W}_{lm}(\omega_{mwv})$ is the Wronskian.
@@
$\omega_{mwv}^{(2)}=m\frac{u^{\varphi}}{u^{t}}+w\frac{\Omega}{u^{t}}+v\frac{\omega_{\mathrm{IB}}}{u^{t}},$ (38)
@@ II.7 Adiabatic evolution of the system
Integrating the $01$ component over the solid angle yields the average energy fluxes at infinity and at the event horizon $r_{+}$. They can be expressed as
@@
where $q_{1}$ and $q_{2}$ are functions of $\mathcal{E}$, $\mathcal{L}_{z}$ and $p$, and these two functions are derived as follows. Since circular orbits remain circular during radiation reaction [97, 98, 99, 100], we have $\dot{R}=0$ and $\dot{R}^{\prime}=0$ according to the radial geodesic equation defined in Eq. (5). The former equation relates the energy and angular-momentum fluxes to the properties of the circular-equatorial orbits, which is
@@ III.1 Frequencies and amplitudes of the modes
$\Omega_{G}\equiv\frac{u^{\varphi}}{u^{t}},\quad\Omega_{P}\equiv\frac{\Omega}{u^{t}},\quad\Omega_{I}\equiv\frac{\omega_{\mathrm{IB}}}{u^{t}}.$ (53)
@@ IV.1 Amplitude, Flux, and resonant excitation of QNMs
Table 2: The zeroth and second order amplitudes and their fractions of main modes, with parameters $a=0.9M$, $p=10M$, $\epsilon=10^{-5}$, $d=4000\mu$, $\tilde{\varphi}=\pi/4$ and $\tilde{\theta}=\pi/3$.
@@
It is remarkable that the ratios $A_{lmwv}$ for the high-frequency modes with $v=\pm 2$ can approach unity in some cases. These second-order relative amplitudes are about $\epsilon^{-1}$ times larger than those of the low-frequency modes. Moreover, for each $m$, there are ten high-frequency modes corresponding to $w=0,\pm 1,\pm 2,\ v=\pm 2$. Therefore, the fluxes are amplified by another order of magnitude. Such a large flux due to the high-frequency modes should cause the b-EMRI system to evolve faster than a standard EMRI. We notice that an earlier study using the MPD equation to solve the evolution of b-EMRI also revealed the appearance of two high-frequency modes (see Eq. (30) in Ref. [67]). However, the authors disregarded these two terms in the calculation of the flux, which potentially could result in a slower evolution of their b-EMRI.
@@
From the table we find that the contributions from the b-EMRI high-frequency modes are substantial, particularly during the early inspiral stage. This leads to a faster evolution of the b-EMRI system compared to a standard EMRI. In contrast, the low-frequency modes contribute only at the $\mathcal{O}(\epsilon)$ order. The behavior of the angular momentum fluxes mirrors that of the energy fluxes and hence are not shown here.
@@
Figure 4: Waveform snapshots with different setting of $d$. Other parameters are fixed at $a=0.9M$, $p=10M$, $\epsilon=10^{-5}$, $\tilde{\varphi}=\pi/4$, $\tilde{\theta}=\pi/3$ with viewing angles $\theta=\pi/3$ and $\varphi=\pi/6$.
@@ IV.2 Adiabatic trajectory and waveform
In this subsection, we study the adiabatic evolution of the b-EMRI system with different configurations and compare it with a standard EMRI. We continue to use the parameters $a=0.9M$, $p=10M$, $\epsilon=10^{-5}$, $d=4000\mu$, $\tilde{\varphi}=\pi/4$, and $\tilde{\theta}=\pi/3$. We also adopt the configurations defined in Eq. (60). Eq. (60a) corresponds to the evolution of a standard EMRI. Eq. (60b) represents the evolution of a b-EMRI driven only by the low-frequency modes, and we refer to this case as “Configuration 1”. Eq. (60c) includes all modes, which we refer to as “Configuration 2”.
@@
These results suggest that the high-frequency modes contribute significantly more to the evolution of the b-EMRI system than the low-frequency modes. As a result, the b-EMRI system evolves much faster than a standard EMRI in the early inspiral phase. This observation extends beyond the current understanding of b-EMRI system evolution. For example, in Fig. 2 of Ref. [67] where the contributions of the high-frequency modes are not included, we see a much slower evolution of the b-EMRI system which is consistent with what we have derived in Fig. 6(a).
@@
Table 4: Energy fluxes due to different modes when $a=0.9M$, $\epsilon=10^{-5}$, $\tilde{\varphi}=\pi/4$, and $\tilde{\theta}=\pi/3$
@@ V.1 Comparisons with previous works
Later, Ref. [67] treats the inner binary as a mass distribution and expands it into multipole moments to compute the source term. To evolve the outer orbit, the authors treat the BBH as a single spinning body and use the MPD equation to compute its orbit around the SMBH. This approach ensures conservation of the stress-energy tensor and inspires us to discuss the validity of our model in Sec. V.3. However, in their study the authors omitted the high-frequency source terms because the frequencies are outside the LISA band. Our results, nevertheless, shows that these dropped high-frequency modes, if resonate with the central SMBH’s QNMs, will enhance energy dissipation. Notably, the IB is allowed to evolve and merger in Ref. [67], so the b-EMRI will be an important multi-band GW source which can produce both high-frequency GW in the LIGO/Virgo band and low-frequency GW in the LISA band [41]. The merger of the IB also leads to a sudden mass loss making the low-frequency waveform more distinguishable. These potential multi-band observational aspects are not discussed in our work.
@@ V.2 Parameter space
Throughout this work, we assume no evolution of the IB, which means that the IB is neither tidally disrupted nor excited to high eccentricities. The stability of a hierarchical three-body system is a classical problem [111]. The maximum distance for the two bodies in the IB to remain stable is determined by the Hill radius, $R_{H}:=(\epsilon/3)^{1/3}p$, where Newtonian gravity is assumed. Since here we are working in the strong-gravity regime, we require $d<\eta\cdot R_{H}$, where $10<\eta<$ 1. The determination of $\eta$ involves more delicate considerations of relativistic three-body dynamics. We refer interested readers to our previous work [77] and a more recent work [60]. Here, we simply set $\eta=1/3$ as an example. Given our fiducial model in which $p=10M$ and $\epsilon=10^{-5}$, we find $d<0.05M$. In other words, an IB with $d=4000\mu=0.04M$ is stable in this case. Additionally, since we assume Newtonian motion for the inner orbital motion, $d$ should be sufficiently large to avoid strong relativistic correction. For this latter reason, we set $d>1000\mu=1000M\epsilon$. The leading post-Newtonian effect to the inner orbit is of the order $\mathcal{O}(\mu\sqrt{\mu/d^{3}})\sim 10^{-6}$, which is negligible. Moreover, the evolution timescale of the IB due to GW radiation is much longer than that of the OB, so we do not the IB here.
Another important assumption in this work is that the inner binary follows a circular orbit. However, dynamical effects in triple systems, such as the Kozai-Lidov mechanism [112, 113], could excite the eccentricity of the IB if the inclination angle $\tilde{\theta}$ between the inner and outer orbits exceeds a critical value [60]. The timescale of the eccentric Kozai-Lidov mechanism (EKM) was derived in Ref. [114, 115], which gives
@@
Figure 7: Parameter space of a b-EMRI. The regions below the solid (dashed) lines satisfy the EKM (Hill’s) condition when $\tilde{p}=20,10,5$. The grey shaded region represents a IB in the mass range of $10M_{\odot}<\mu<100M_{\odot}$, where we have assumed $M=4\times 10^{6}M_{\odot}$. The red star denotes our fiducial model in which $\epsilon=10^{-5},\ \tilde{d}=4000$.
@@ V.3 Validity of the model
In Sec. II.7 we demonstrated how to track the evolution of the system ‘quasi-adiabatically’. In particular, the evolution of the parameter $p$ is derived assuming osculating geodesic motion of the OB. In this work, however, we follow Ref. [65, 66] and write the stress-energy tensor of the IB as
@@
While we have not yet included this interaction term in the stress-energy when solving the Teukolsky equation, we remark that it is straight forward to add in follow up work as a linear correction to the Teukolsky amplitudes (and therefore fluxes). Without accounting for that term, the waveform model should not be used in parameter estimation studies for example. However, we stress that its neglect does not change the qualitative features of our findings.
@@ Acknowledgements.
This work is supported by the National Key Research and Development Program of China (Grant No. 2021YFC2203002) and the National Natural Science Foundation of China (Grant No. 12473037). We would like to acknowledge our referee for thoroughly reading the manuscript and providing constructive suggestions. Y.Y. acknowledges NUS Physics Department for providing office environment and care. Y.Y. acknowledges Haolan Zheng and Han Yan for several inspirational discussions. The authors thank Soichiro Isoyama, Wenbiao Han, and Xilong Fan for insightful feedback on a draft of this paper. J.M. acknowledges support by the NUS Faculty of Science, under the research grant 22-5478-A0001.
@@ Appendix A Transformation from the LIF to BL coordinates
$c_{01}=$ $\frac{\mathcal{M}\mathcal{M}_{\varphi}+\gamma_{\varphi\varphi}^{1/2}\nu}{4\gamma_{\varphi\varphi}^{1/2}\mathcal{M}/\gamma}\Big{\{}(1+\cos\tilde{\theta})\sin\left[\left(\Omega+\omega_{\mathrm{IB}}\right)\tau-\tilde{\varphi}\right]$
$-(1-\cos\tilde{\theta})\sin\left[\left(\Omega-\omega_{\mathrm{IB}}\right)\tau-\tilde{\varphi}\right]\Big{\}},$ (72a)
@@
$c_{11}=$ $\frac{1}{4\gamma_{rr}^{1/2}}\Big{\{}(1+\cos\tilde{\theta})\cos\left[\left(\Omega+\omega_{\mathrm{IB}}\right)\tau-\tilde{\varphi}\right]$
$-(1-\cos\tilde{\theta})\cos\left[\left(\Omega-\omega_{\mathrm{IB}}\right)\tau-\tilde{\varphi}\right]\Big{\}},$ (72b)
@@
$c_{21}=$ $\frac{1}{2\gamma_{\theta\theta}^{1/2}}\sin\tilde{\theta}\cos\left(\omega_{\mathrm{IB}}\tau\right),$ (72c)
@@
$c_{31}=$ $\frac{\gamma}{4\gamma_{\varphi\varphi}^{1/2}}\Big{\{}(1+\cos\tilde{\theta})\sin\left[\left(\Omega+\omega_{\mathrm{IB}}\right)\tau-\tilde{\varphi}\right]$
$-(1-\cos\tilde{\theta})\sin\left[\left(\Omega-\omega_{\mathrm{IB}}\right)\tau-\tilde{\varphi}\right]\Big{\}},$ (72d)
$c_{02}=$ $\frac{(\mathcal{M}\mathcal{M}_{\varphi}+\gamma_{\varphi\varphi}^{1/2}\nu)\kappa_{\varphi}-(\gamma_{\varphi\varphi}^{1/2}+\mathcal{M}\mathcal{M}_{\varphi}\nu)\Omega}{16\gamma_{\varphi\varphi}^{1/2}\mathcal{M}/\gamma}$
@@
×$\times\Big{\{}2\cos^{4}\frac{\tilde{\theta}}{2}\sin\left[2\left(\Omega+\omega_{\mathrm{IB}}\right)\tau-2\tilde{\varphi}\right]$
$-\sin^{2}\tilde{\theta}\sin\left(2\Omega\tau-2\tilde{\varphi}\right)$
@@
$+2\sin^{4}\frac{\tilde{\theta}}{2}\sin\left[2\left(\Omega-\omega_{\mathrm{IB}}\right)\tau-2\tilde{\varphi}\right]\Big{\}},$ (72e)
$c_{12}=$ $\frac{\kappa_{r}+\kappa_{\varphi}}{16\gamma_{rr}^{1/2}}\cos^{4}\frac{\tilde{\theta}}{4}\cos\left[2\left(\Omega+\omega_{\mathrm{IB}}\right)\tau-2\tilde{\varphi}\right]$
$+\frac{\kappa_{r}+\kappa_{\varphi}}{16\gamma_{rr}^{1/2}}\sin^{4}\frac{\tilde{\theta}}{2}\cos\left[2\left(\Omega-\omega_{\mathrm{IB}}\right)\tau-2\tilde{\varphi}\right]$
$-\frac{\kappa_{r}+\kappa_{\varphi}}{32\gamma_{rr}^{1/2}}\sin^{2}\tilde{\theta}\cos\left(2\Omega\tau-2\tilde{\varphi}\right)$
$-\frac{\kappa_{r}+2\kappa_{\theta}-\kappa_{\varphi}}{32\gamma_{rr}^{1/2}}\sin^{2}\tilde{\theta}\cos\left(2\omega_{\mathrm{IB}}\tau\right)$
$+\frac{\kappa_{r}-\kappa_{\varphi}}{32\gamma_{rr}^{1/2}}\left(1+\cos^{2}\tilde{\theta}\right)-\frac{\kappa_{\theta}}{16\gamma_{rr}^{1/2}}\sin^{2}\tilde{\theta},$ (72f)
$c_{22}=$ $\frac{\kappa_{\theta}\sin\tilde{\theta}}{16\gamma_{\theta\theta}^{1/2}}\Big{\{}\left(1+\cos\tilde{\theta}\right)\cos\left[\left(\Omega+2\omega_{\mathrm{IB}}\right)\tau-\tilde{\varphi}\right]$
$-\left(1-\cos\tilde{\theta}\right)\cos\left[\left(\Omega-2\omega_{\mathrm{IB}}\right)\tau-\tilde{\varphi}\right]$
$+2\cos\tilde{\theta}\cos\left(\Omega\tau-\tilde{\varphi}\right)\Big{\}},$ (72g)
$c_{32}=$ $\frac{\kappa_{\varphi}-\nu\Omega}{16\gamma_{\varphi\varphi}^{1/2}/\gamma}\times\Big{\{}2\cos^{4}\frac{\tilde{\theta}}{2}\sin\left[2\left(\Omega+\omega_{\mathrm{IB}}\right)\tau-2\tilde{\varphi}\right]$
$-\sin^{2}\tilde{\theta}\sin\left(2\Omega\tau-2\tilde{\varphi}\right)$
$+2\sin^{4}\frac{\tilde{\theta}}{2}\sin\left[2\left(\Omega-\omega_{\mathrm{IB}}\right)\tau-2\tilde{\varphi}\right]\Big{\}}.$ (72h)
As we can see, these coefficients are all functions of $\tau$.